# A Bilevel Optimization Framework for Imbalanced Data Classification

### v1

@@ 1 Introduction
Technological advances have ushered in an era where data generated before, during, and after a scientific experiment is kept for further analysis. Such vast quantities of stored data have changed the landscape of analysis where careful attention must be paid to new statistical challenges introduced by massive amounts of data [1]. For instance, a billion subatomic particle interactions take place over a period of 24 hours in the ATLAS particle detector. However, only about 2,500 of these interactions are useful and must be extracted through careful analysis. This problem of identifying a small amount of useful events from a large pool of interactions is often modelled as an imbalanced data classification problem [2].
@@
A dataset is imbalanced when one of the categories represented in the data is strongly under-represented. As depicted in Figure 1(a), the under-represented class is called minority data whereas the over-represented class is called majority data. Other examples of imbalanced data sets include diseases [3, 4, 5]; chemicals either harmful or beneficial to humans [6, 7]; natural disasters [8]; fraudulent transactions [9, 10] and many more [11, 12]. While it is the rare minority data that is of interest to identify and analyze, the sparsity of the minority data within the larger dataset presents challenges to neural network models.
A fundamental problem of classifying imbalanced data is how the imbalance impairs training. Without intervention and as depicted in Figure 1(b), training sets of imbalanced data are also imbalanced. Without an equal amount of minority and majority data for training, a model is impaired in its ability to classify the one it lacks - the minority data.
@@
To reduce the issues due to noise and overlap, a tool originally built to generate realistic images called Generative Adversarial Networks (GANs)[14] has been repurposed. Specifically, GANs have been repurposed to leverage their adversarial structure to reject synthetically generated datapoints that do not help with classification [15, 16, 17]. More specifically, GANs reject synthetic datapoints that do not decrease classification loss. Despite improvements made with the use of GANs, problems of noise and overlap persist when using oversampling methods to classify imbalanced data. While GANs help with eliminating synthetic datapoints that do not improve model loss, they do not necessarily eliminate synthetic datapoints contributing to noise and overlap. Given that the number of different features making up a single datapoint can be quite large, and that each individual feature can also have a large number of possible values, there is a vast sea of possibilities for any one datapoint. Therefore, creating synthetic data that matches minority data in a way that distinguishes it from majority data while also avoiding problems of over-fitting is a challenge difficult to overcome.
@@ 2 Background
To measure a classifier’s performance, let $T_{p}\in\mathbb{N}$ and $F_{n}\in\mathbb{N}$ denote the number of correctly (truly) and incorrectly (falsely) classified minority datapoints, respectively; which implies $m=T_{p}+F_{n}$. Similarly, let $T_{n}\in\mathbb{N}$ and $F_{p}\in\mathbb{N}$ denote the number of correctly (truly) and incorrectly (falsely) classified majority datapoints; $M=T_{n}+F_{p}$. w.l.o.g., let $\hat{y}_{1},\dots,\hat{y}_{m}$ be the predicted labels of minority datapoints and $\hat{y}_{m+1},\dots,\hat{y}_{m+M}$ be the predicted labels of majority datapoints. Therefore,
$T_{p}=\sum_{i=1}^{m}\hat{y}_{i}$
$F_{n}=m-T_{p}$
$F_{p}=\sum_{i=m+1}^{m+M}\hat{y}_{i}$
and
$T_{n}=M-F_{p}.$
For example, let this row of $1$’s and 00’s be a vector of actual labels; i.e., $y\in\{0,1\}^{10}$, with the labels of three minority datapoints on the left and the labels of seven majority datapoints on the right.
11111111111100000000000000
An example classification of the datapoints associated with the 10 labels above is this vector of predicted labels, i.e., $\hat{y}\in\{0,1\}^{10}$, below. Counting the total numbers of correctly and incorrectly classified minority datapoints on the left ($t_{p}$/$f_{n}$), $T_{p}=2$ and $F_{n}=1$; and of correctly and incorrectly classified majority datapoints on the right ($t_{n}$/$f_{p}$), $T_{n}=5$, and $F_{p}=2$.
[…]
@@
Definition 3.
Let $l(s;z(w))\in\mathbb{R}^{+}$ denote loss as follows
$l(s;z(w))$ =$=\begin{cases}-\log(p(y=0|x)),\quad\text{if }s\in S_{M}\\ -\log(p(y=1|x)),\quad\text{if }s\in S_{m}\\ \end{cases}$
=$=\begin{cases}-\log(\frac{e^{1-z(w)}}{e^{z(w)}+e^{1-z(w)}}),\quad\text{if }s\in S_{M}\\ -\log(\frac{e^{z(w)}}{e^{z(w)}+e^{1-z(w)}}),\quad\text{if }s\in S_{m}\\ \end{cases}$
Definition 4.
Let $\mathscr{L}(S;z(w))$ denote sample loss as follows
$\mathscr{L}(S;z(w))$ $=\frac{1}{|S|}\sum_{i=1}^{|S|}l(s_{i};z(w))$
@@ 3 Bilevel Optimization Formulation
Figure 2: MUBO’s bilevel optimization formulation features a lower level problem whose solution is optimal model parameters and an upper level problem that identifies an optimal collection of majority data subsamples. The outer loop accepts only subsamples that decrease majority loss.
@@
where $S=S_{m}\cup S_{M}$.
Depicted in Figure 2, the upper level problem choses an optimal majority subsample, $S_{M}$, to minimize the majority loss function, subject to finding optimal neural network weights to minimize the sample loss $\mathscr{L}(S;z(w))$. Bilevel optimization is also referred to as leader-follower games and is a sequential decision model where the follower (lower level problem) chooses optimal weights given the leader’s (upper level) choice of sample set. In this sense, the leader can anticipate the follower’s decision and takes the optimal weights into account when choosing the sample set $S_{M}$.
Our unique use of bilevel optimization resides in the formulation of the upper level problem. By focusing the upper level problem solely on finding an optimal subset of majority training data, we increase the likelihood of correctly classifying majority data. Additionally, using optimal majority training data, in turn, also increases the likelihood of correctly classifying minority data as there are only two classes. In turn, the opportunity for overlap between majority and minority data decreases. Attaining a noise-less and balanced subset of training data, we pair the resulting optimal subset of majority data with the full set of minority training data.
To note, a bilevel optimization framework lends itself well to other classification methods and learning more generally. The upper level problem, in particular, can be swapped out for achieving other learning objectives.
@@ 7       $S_{t}=S_{t-1}\cup S_{M_{t}}$
8       $J(S_{M_{t}})=\text{innerLoop(}S_{t})-\text{ innerLoop(}S_{t})[1:m]$
9       $\gamma_{t}=J(S_{M_{t-1}})$
@@ 12             $J(S_{M_{t}})=J(S_{M_{t-1}})$
13
14      else if $|S_{t-1}|\leq\frac{7}{4}m$ then
15            $M_{t}=\lceil\frac{1}{4}m\rceil$
@@
17                  $M_{t}=\lceil\frac{1}{100}m\rceil$
18
@@ 19      $t=t+1$
21function $\text{innerLoop}{(}S_{t}{)}$
@@ 24             $epochs=epochs+1$
25
26      return $\mathscr{L}$
27
28function $\text{checkStep}{(}J{(}S_{M_{t}}{)},\gamma_{t}{)}$
29       if $J(S_{M_{t}})>\gamma_{t}$ then
@@ 4.1 Motivation for the Upper Level Problem
To begin, we recall from Definitions 1 and 3 the connection between a majority datapoint’s loss $l(s;z(w))$ and whether or not its predicted label $\hat{y}$ is correct or incorrect. With the following lemma, we extend the connection between loss and classification to show that loss of a correctly classified majority datapoint is strictly less than the loss of an incorrectly classified majority datapoint.
@@ 5 Evaluation
Accuracy 0.98 0.98 0.98 0.98 0.99 0.99 0.99 0.99
Table 1: Average $F1$, precision, recall and accuracy scores provide a comparison of MUBO to a technique without sampling (w/o); three classic sampling techniques - SMOTE, Borderline-SMOTE (B-SMO) and SVM-SMOTE (S-SMO); and three state-of-the-art methods - SMOTified-GAN (SMO-GAN), GBO and SSG. Experiments across all methods are with $20\%$ testing data.
@@ 5.2 Metrics
In addition to the F1 scores previously defined in Definition 2, the following evaluation metrics are used:
@@
Both precision and recall are important for evaluating imbalanced data classification. To differentiate between precision and recall, imagine two bins - one bin for minority data and one bin for majority. Low precision indicates that data has been put into the wrong bin. Low recall, on the other hand, indicates that a bin is missing some of its own datapoints. For example, high precision is important for fraud detection so as to avoid the false accusation of fraud. High recall is important for medical image classification as missing a diagnosis can have fatal consequences. As the harmonic mean between precision and recall, the F1 score captures both precision and recall and is the preferred metric for evaluating imbalanced data classification. [41]
@@ 5.3 Numerical results and analysis
SSG 0.95 0.90 1.00 0.91 0.83 1.00 0.98 0.99 0.99
MUBO 0.99 0.99 0.98 0.99 0.97 1.00 0.99 1.00 0.97
Table 3: $F1$ (average), $F1_{m}$ (minority), $F1_{M}$ (majority), $Prec.$ (average precision), $Prec_{m}$ (minority), $Prec_{M}$ (majority), $Recall$ (average recall), $Rec_{m}$ (minority) and $Rec_{M}$ (majority) scores provide a comparison of minority classification performance versus majority classification performance for MUBO and two state-of-the-art methods GBO and SSG.
As shown in Table 1, the average F1 scores for our method surpass the average F1 scores of all classic and state-of-the-art sampling methods on the datasets Gisette, Abalone and Connect4. MUBO’s leading F1 scores range from $0.86$ to $0.99$ (with variances from $1.31e{-5}$ to $2.29e{-5}$), and the next highest scores are $2\%$ to $4\%$ lower. The two datasets for which our approach does not lead – Spambase and Ionosphere – are, in fact, the least imbalanced. In comparison with state-of-the-art methods, MUBO met SMOTified-GAN’s F1 score with the Spambase dataset while achieving an F1 score just shy of that of GBO and SSG at only $1\%$ lower. On the Ionosphere dataset, the average F1 score of our method is $11\%$ lower than that of SMOTified-GAN and $5\%$ lower than GBO and SSG. A deeper investigation into these results as shown in Table 3 reveal that Ionosphere’s majority F1 score is $0.93$ while its minority F1 score is $0.81$, indicating that our technique struggled to classify minority data. Ionosphere is the smallest dataset in the study with 351 instances. Given that our method utilizes all available minority data, Ionosphere’s experimental results indicate that a small dataset size is a hindrance to our undersampling approach.
Table 3 provides a comparison between minority and majority classification performance. On all five datasets, $F1_{m}<F1_{M}$ for GBO and SSG, signalling that these minority oversampling methods reliant on synthetic data are better at classifying majority data than minority data. In contrast, our technique does not use synthetic data and, therefore, sidesteps potential hurdles to minority data classification. Aside from Ionosphere, our method’s minority F1 score surpasses the minority F1 score for GBO and SSG, with the next highest score $1\%$ to $19\%$ lower. Our approach exhibits less optimal performance on majority recall. Given that the majority subset with which MUBO trains is significantly smaller than the full set of majority data with which GBO and SSG train, a lower majority recall score makes sense.
@@ 6 Conclusion and Future Outlook
Additional conditions are required to formally prove experimental results showing improved majority loss implies improved average F1 score. For starters, we need a bound on minority loss $J(S_{m})$. Recall from Section 4 that loss of truly classified majority data is strictly less than loss of falsely classified majority data; i.e., $J_{M}^{T}<J_{M}^{F}$. Additional conditions required to prove theoretical results include bounds on loss of falsely classified minority datapoints and bounds on the differences between loss associated with truly and falsely classified datapoints. Future work includes the formulation of subsequent methods that take into account these additional conditions for proving improved results.

### v2

@@ 1 Introduction
Technological advances have ushered in an era when all the data generated for before, during, and after a scientific experiment is kept. Such vast quantities of stored data have changed the landscape of analysis where careful attention must be paid to new statistical challenges introduced by massive amounts of data [1]. For instance, a billion subatomic particle interactions take place over a period of 24 hours in the ATLAS particle detector. However, only about 2,500 of these interactions are useful and must be extracted through careful analysis. This problem of identifying a small amount of useful events from a large pool of interactions is often modelled as an imbalanced data classification problem [2].
@@
A dataset is imbalanced when one of the categories represented in the data is strongly under-represented. As depicted in Figure 1(a), the under-represented class is called minority data whereas the over-represented class is called majority data. Other examples of imbalanced data sets include diseases [3, 4, 5]; chemicals either harmful or beneficial to humans [6, 7]; natural disasters [8]; fraudulent transactions [9, 10] and many more [11, 12]. While it is the rare minority data that is of interest, the sparsity of the minority data within the larger dataset presents challenges to neural network models.
A fundamental problem of classification is the impairment that imbalanced data causes to training. Without intervention and as depicted in Figure 1(b), training sets of imbalanced data are also imbalanced. Without an equal amount of minority and majority data for training, a model is impaired in its ability to classify the one it lacks - the minority data.
@@
To reduce the issues due to noise and overlap, a tool originally built to generate realistic images called Generative Adversarial Networks (GANs)[14] has been repurposed. Specifically, GANs have been repurposed to leverage their adversarial structure to reject synthetically generated datapoints that do not help with classification [15, 16, 17]. More specifically, GANs reject synthetic datapoints that do not decrease classification loss. Despite improvements made with the use of GANs, problems of noise and overlap persist when using oversampling methods to classify imbalanced data.
While GANs help with eliminating synthetic datapoints that do not improve model loss, they do not necessarily eliminate synthetic datapoints contributing to noise and overlap. Given that the number of different features making up a single datapoint can be quite large, and that each individual feature can also have a large number of possible values, there is a vast sea of possibilities for any one datapoint. Therefore, creating synthetic data that matches minority data in a way that distinguishes it from majority data while also avoiding problems of over-fitting is a challenge difficult to overcome.
@@ 2 Background
Definition 2.
Let $l(s;z(w))\in\mathbb{R}^{+}$ denote loss as follows
$l(s;z(w))$ =$=\begin{cases}-\log(\frac{e^{1-z(w)}}{e^{z(w)}+e^{1-z(w)}}),\quad\text{if }s\in S_{M};\text{ i.e.},y=0\\ -\log(\frac{e^{z(w)}}{e^{z(w)}+e^{1-z(w)}}),\quad\text{if }s\in S_{m};\text{ i.e.},y=1\\ \end{cases}$
Definition 3.
Let $\mathscr{L}(S;z(w))$ denote sample loss as follows
$\mathscr{L}(S;z(w))$ $=\frac{1}{|S|}\sum_{i=1}^{|S|}l(s_{i};z(w))$
To measure a classifier’s performance, let $T_{p}\in\mathbb{N}$ and $F_{n}\in\mathbb{N}$ denote the number of correctly (truly) and incorrectly (falsely) classified minority datapoints, respectively; which implies $m=T_{p}+F_{n}$. Similarly, let $T_{n}\in\mathbb{N}$ and $F_{p}\in\mathbb{N}$ denote the number of correctly (truly) and incorrectly (falsely) classified majority datapoints; $M=T_{n}+F_{p}$. w.l.o.g., let $\hat{y}_{1},\dots,\hat{y}_{m}$ be the predicted labels of minority datapoints and $\hat{y}_{m+1},\dots,\hat{y}_{m+M}$ be the predicted labels of majority datapoints. Therefore, $T_{p}=\sum_{i=1}^{m}\hat{y}_{i}$, $F_{n}=m-T_{p}$, $F_{p}=\sum_{i=m+1}^{m+M}\hat{y}_{i}$ and $T_{n}=M-F_{p}$.
@@ 3 Bilevel Optimization Formulation
where $S=S_{m}\cup S_{M}$. Figure 2 depicts the connection between MUBO and its bilevel problem.
Figure 2: MUBO is designed to solve its bilevel optimization problem. One full iteration is shown above with the following steps: i. micro sample majority training data (the number of datapoints in the micro-sample is significantly less than the number of minority datapoints); ii. train the model using the new majority micro-sample combined with the previously accepted micro-samples of majority data and all minority training data, $S_{t}=S_{M_{t}}\cup S_{m}$, finding optimal weights, $w^{*}_{t}$, that minimize overall loss (inner optimization step); iii. compute majority loss using the optimal weights, $J(S_{M_{t}};w_{t}^{*})=\mathscr{L}(S_{t};w_{t}^{*})-J(S_{m};w_{t}^{*})$; iv. compare majority loss of the current step $t$ with that of step $t-1$, accepting the new micro-sample of majority training data only if it improves majority loss, i.e., $J(S_{M_{t}};w_{t}^{*})\leq J(S_{M_{t-1}};w_{t-1}^{*})$; and v. repeat steps i-iv to find a subset of majority training data that minimizes majority loss (outer optimization step).
The outer level problem chooses an optimal majority subsample, $S_{M}$, to minimize the majority loss function, subject to finding optimal neural network weights to minimize the sample loss $\mathscr{L}(S;z(w))$. Bilevel optimization is also referred to as leader-follower games and is a sequential decision model where the follower (inner level problem) chooses optimal weights given the leader’s (outer level) choice of sample set. In this sense, the leader can anticipate the follower’s decision and takes the optimal weights into account when choosing the sample set $S_{M}$.
Our unique use of bilevel optimization resides in the formulation of the outer level problem. By focusing the outer level problem solely on finding an optimal subset of majority training data, we increase the likelihood of correctly classifying majority data. Additionally, using optimal majority training data, in turn, also increases the likelihood of correctly classifying minority data as there are only two classes. In turn, the opportunity for overlap between majority and minority data decreases. Attaining a noise-less and balanced subset of training data, we pair the resulting optimal subset of majority data with the full set of minority training data.
To note, a bilevel optimization framework lends itself well to other classification methods and learning more generally. The outer level problem, in particular, can be swapped out for achieving other learning objectives.
@@ 7       $S_{t}=S_{t-1}\cup S_{M_{t}}$
8       $J(S_{M_{t}})=\text{mean.innerLoop(}S_{t})-\text{ mean.innerLoop(}S_{t})[1:m]$
@@ 11             $J(S_{M_{t}})=J(S_{M_{t-1}})$
12
13      if $|S_{t-1}|\leq\frac{7}{4}m$ then
14            $M_{t+1}=\lceil\frac{1}{10}m\rceil$
@@
16                  $M_{t+1}=\lceil\frac{1}{100}m\rceil$
17
@@ 18      $t=t+1$
20function $\text{innerLoop}{(}S_{t}{)}$
@@ 23             $epochs=epochs+1$
24
25      return $\mathbf{l(s)}[1:m+M_{t}]$
26
27function $\text{checkStep}{(}J{(}S_{M_{t}}{)},J{(}S_{M_{t-1}}{)}{)}$
28       if $J(S_{M_{t}})>J(S_{M_{t-1}})$ then
@@ 4.1 Motivation for the Upper Level Problem
To begin, we recall from Definitions 1 and 2 the connection between a majority datapoint’s loss $l(s;z(w))$ and whether or not its predicted label $\hat{y}$ is correct or incorrect. With the following lemma, we extend the connection between loss and classification to show that loss of a correctly classified majority datapoint is strictly less than the loss of an incorrectly classified majority datapoint.
@@ 5 Evaluation
Table 1: Average $F1$, precision and recall scores provide a comparison of MUBO to a technique without sampling (w/o); three classic sampling techniques - SMOTE, Borderline-SMOTE (B-SMO) and SVM-SMOTE (S-SMO); and three state-of-the-art methods - SMOTified-GAN (SMO-GAN), GBO and SSG. Experiments across all methods are with $20\%$ testing data.
@@ 5.2 Metrics
In addition to the F1 scores previously defined in Definition 4, the additional two metrics are used:
@@
Both precision and recall are important for evaluating imbalanced data classification. To differentiate between precision and recall, imagine two bins - one bin for minority data and one bin for majority. Low precision indicates that data has been put into the wrong bin. Low recall, on the other hand, indicates that a bin is missing some of its own datapoints. For example, high precision is important for fraud detection so as to avoid the false accusation of fraud. High recall, on the other hand, is important for medical image classification as missing a diagnosis can have fatal consequences. As the harmonic mean between precision and recall, the F1 score captures both precision and recall and is the preferred metric for evaluating imbalanced data classification. [41]
@@ 5.3 Numerical results and analysis
MUBO 0.96 0.91 1.00 0.92 0.85 1.00 0.99 0.99 0.99
Table 3: A comparison between minority classification versus majority classification, $F1$ (average), $F1_{m}$ (minority), $F1_{M}$ (majority), $Prec.$ (average precision), $Prec_{m}$ (minority), $Prec_{M}$ (majority), $Recall$ (average recall), $Rec_{m}$ (minority) and $Rec_{M}$ (majority) scores are provided for MUBO and the state-of-the-art method GBO.
As shown in Table 1, the average F1 scores for our method surpass the average F1 scores of all classic and state-of-the-art sampling methods on the datasets Gisette, Abalone and Connect4. MUBO’s leading F1 scores range from $0.86$ to $0.96$ (with variances from $1.31e{-5}$ to $2.29e{-5}$), and the next highest scores are $1\%$ to $2\%$ lower. The two datasets for which our approach does not lead – Spambase and Ionosphere – are, in fact, the least imbalanced. In comparison with state-of-the-art methods, MUBO met SMOTified-GAN’s F1 score with the Spambase dataset while achieving an F1 score just shy of that of GBO and SSG at only $1\%$ lower. On the Ionosphere dataset, the average F1 score of our method is $11\%$ lower than that of SMOTified-GAN and $5\%$ lower than GBO. A deeper investigation into these results as shown in Table 3 reveal that Ionosphere’s majority F1 score is $0.93$ while its minority F1 score is $0.81$, indicating that our technique struggled to classify minority data. Ionosphere is the smallest dataset in the study with 351 instances. Given that our method utilizes all available minority data, Ionosphere’s experimental results indicate that a small dataset size is a hindrance to our undersampling approach.
Table 3 provides a comparison between minority and majority classification performance. On all five datasets, $F1_{m}<F1_{M}$ for GBO, signaling that this minority oversampling method reliant on synthetic data is better at classifying majority data than minority data. In contrast, our technique does not use synthetic data and, therefore, sidesteps potential hurdles to minority data classification. Aside from Ionosphere, our method’s minority F1 score surpasses the minority F1 score for GBO, with the next highest score $1\%$ to $19\%$ lower. Our approach exhibits less optimal performance on majority recall. Given that the majority subset with which MUBO trains is significantly smaller than the full set of majority data with which GBO trains, a lower majority recall score makes sense.
@@ 6 Conclusion and Future Outlook
Additional conditions are required to formally prove experimental results showing improved majority loss implies improved average F1 score. Future work includes the formulation of subsequent methods that provide conditions for proving improved results. To add, the formulation of subsequent methods also includes a generalization of current work to include strong results on both small and large datasets.